rectangular room has width half of its length. When 6 is decreased from both length and width then its area is differed by 108 so find the width.
Answer by mananth(16946)

let width be x
length will be 2x
..
6 is decreased from both
x-6
2x-6
2x^2-(x-6)(2x-6)= 108
2x^2-(2x^2-6x-12x+36)=108
2x^2-2x^2+18x-36=108
18x=108+36
18x=144
x=144/18
x=8 which is the width
Length = 2x = 16 units